OpenStudy (lukecrayonz):

2.4e^(x-6)=9.3 2.4e^(x-6)=9.3 @Mathematics

OpenStudy (anonymous):

around 7.35

OpenStudy (anonymous):

e^(x-6)=9.3/2.4=3.875 ln(3.875)=x-6 1.35=x-6 (approximately) x=7.35